# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v3

@@ Abstract
We analyse the relationship between privacy vulnerability and dataset properties, such as examples per class and number of classes, when applying two state-of-the-art membership inference attacks (MIAs) to fine-tuned neural networks. We derive per-example MIA vulnerability in terms of score distributions and statistics computed from shadow models. We introduce a simplified model of membership inference and prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class. We complement the theoretical analysis with empirical analysis by systematically testing the practical privacy vulnerability of fine-tuning large image classification models and obtain the previously derived power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. Finally, we fit a parametric model of the previously derived form to predict true positive rate based on dataset properties and observe good fit for MIA vulnerability on unseen fine-tuning scenarios.
@@ 1 Introduction
Machine learning models are prone to memorising their training data, which makes them vulnerable to privacy attacks such as membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and reconstruction attacks (e.g. Balle et al., 2022; Nasr et al., 2023). Differential privacy (DP; Dwork et al., 2006) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications. In health, for example, many European countries and soon also the EU within the European Health Data Space have requirements that models trained on health data that are made publicly available must be anonymous, i.e. they must not contain information that can be linked to an identifiable individual. On the other hand, loss of utility of the model due to privacy constraints may compromise the health benefits that might be gained from it.
In this paper, our aim is to theoretically understand and apply two state-of-the-art MIAs systematically, to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections. We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Feldman & Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning.
In this paper, we theoretically study the relationship between examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning and derive a power-law relationship in the form
@@
We complement the theoretical analysis with extensive experiments over many datasets with varying sizes in the transfer learning setting for image classification tasks and observe the same power law.
Previously, Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how dataset properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by a more systematic study and modelling of dataset properties.
List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art MIAs LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024). Both MIAs are based on the assumption that the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack. We first analyze both MIAs theoretically by introducing a simplified model of membership inference (Section 3). We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models. We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive dataset, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain dataset properties (Section 4).
@@
    Per-example vulnerability: We derive per-example LiRA and RMIA vulnerability (tpr at fixed fpr) in terms of MIA attack score distributions and statistics computed from shadow models (see Section 3.3).
@@
    Power-law in simplified model of MI: We formulate a simplified model of membership inference to quantitatively relate dataset properties and MIA vulnerability. For this model, we prove a power-law relationship between the per-example LiRA and RMIA vulnerability and the number of examples per class. We then extend the the power-law relationship to the average-case LiRA and RMIA vulnerability (See Section 3.4).
@@
    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
  4. 4.
    Regression model: We utilise our empirical observations to fit a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$), which follows the functional form of the theoretically derived power-law. We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 4).
@@ 2 Background
Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model (Shokri et al., 2017). Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
@@
LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not. For a target sample $(x,y_{x})$, LiRA trains the shadow models: 1 with $(x,y_{x})$ as a part of the training set ($(x,y_{x})\in\mathcal{D}$) and 2 without $x$ in the training set ($(x,y_{x})\notin\mathcal{D}$) . After training the shadow models, $(x,y_{x})$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $(x,y_{x})\in\mathcal{D}$ shadow models, and one for the $(x,y_{x})\notin\mathcal{D}$. Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $(x,y_{x})\in\mathcal{D}$ or $(x,y_{x})\notin\mathcal{D}$. We use an optimization by Carlini et al. (2022) for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models. It relies on sampling the shadow datasets in a way that each sample is in expectation half of the time included in the training dataset of a shadow model and half of the time not. At attack time each model will be attacked once using all other models as shadow models.
Robust Membership Inference Attack (RMIA; Zarifzadeh et al., 2024) Recently Zarifzadeh et al. (2024) proposed a new MIA algorithm called RMIA, which aims to improve performance when the number of shadow models is limited. Similar to LiRA, RMIA is based on shadow model training and computing the attack statistics based on a likelihood ratio. The main difference to LiRA is that RMIA does not compute the likelihood ratio based on aggregated IN/OUT statistics, but instead compares the target data point against random samples $(z,y_{z})$ from the target data distribution. After computing the likelihood ratios over multiple $(z,y_{z})$ values, the MIA score is estimated as a proportion of the ratios exceeding a preset bound. This approach makes RMIA a more effective attack when the number of shadow models is low.
@@ 3 Theoretical analysis
In this section, we seek to theoretically understand the impact of the dataset properties on the MIA vulnerability. It is known that different data points exhibit different levels of MIA vulnerability depending on the underlying distribution (e.g.  Aerni et al., 2024; Leemann et al., 2024). Therefore, we start with analysing per-example vulnerabilities for LiRA and RMIA. In order to quantitatively relate dataset properties to these vulnerabilities, a simplified model is formulated. Within this model, we prove a power-law between the per-example vulnerability and the number $S$ of examples per class. Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section 4.
@@ 3.1 Preliminaries
First, let us restate the MIA score from LiRA as defined by Carlini et al. (2022). Denoting the logit of a target model $\mathcal{M}$ applied on a target data point $(x,y_{x})$ as $\ell(\mathcal{M}(x),y_{x})$, the LiRA computes the MIA score as the likelihood ratio
$\mathrm{LR}(x)=\frac{p(\ell(\mathcal{M}(x),y_{x})\mid\mathbb{Q}_{\text{in}}(x,y_{x}))}{p(\ell(\mathcal{M}(x),y_{x})\mid\mathbb{Q}_{\text{out}}(x,y_{x}))},$ (2)
@@
where the $\mathbb{Q}_{\text{in}/\text{out}}$ denote the hypotheses that $(x,y_{x})$ was or was not in the training set of $\mathcal{M}$. Carlini et al. (2022) approximate the IN/OUT hypotheses as normal distributions. Denoting $t_{x}=\ell(\mathcal{M}(x),y_{x})$, the score becomes
$\mathrm{LR}(x)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}^{2}(x))}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}^{2}(x))},$ (3)
@@
where the $\hat{\mu}_{\text{in/out}}(x)$ and $\hat{\sigma}_{\text{in/out}}(x)$ are the means and standard deviations for the IN/OUT shadow model losses for $(x,y_{x})$. Larger values of $\mathrm{LR}(x)$ suggest that $(x,y_{x})$ is more likely in the training set and vice versa. Now, to build a classifier from this score, the LiRA tests if $\mathrm{LR}(x)>\beta$ for some threshold $\beta$.
Next, let us restate how RMIA (Zarifzadeh et al., 2024) builds the MIA score. RMIA augments the likelihood-ratio with a sample from the target data distribution to calibrate how likely you would obtain the target model if $(x,y_{x})$ is replaced with another sample $(z,y_{z})$. Denoting the target model parameters with $\theta$, RMIA computes
$\mathrm{LR}(x,z)=\frac{p(\theta\mid x,y_{x})}{p(\theta\mid z,y_{z})},$ (4)
and the corresponding MIA score is given as
$\text{Score}_{\text{RMIA}}(x)=\Pr_{(z,y_{z})\sim\mathbb{D}}(\mathrm{LR}(x,z)>\gamma),$ (5)
where $\mathbb{D}$ denotes the training data distribution. Similar to LiRA, the classifier is built by checking if $\text{Score}_{\text{RMIA}}(x)>\beta$. In the following, we will use the direct computation of likelihood-ratio as described in Equation 11 of Zarifzadeh et al. (2024) which approximates $\mathrm{LR}(x,z)$ using normal distributions.
3.2 Computing the tpr for LiRA and RMIA
Using the LiRA formulation of Equation 3, the tpr for the target point $(x,y_{x})$ for LiRA is defined as
@@ 3.2 Computing the tpr for LiRA and RMIA
$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (6)
@@
where $\beta$ is a threshold that defines a rejection region of the likelihood ratio test, $\hat{\mu}_{\mathrm{in}}(x),\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{in}}(x)$ and $\hat{\sigma}_{\mathrm{out}}(x)$ are LiRA statistics computed from shadow models, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Appendix A for derivation).
For theoretical analysis of RMIA, we focus on the direct approach that is an approximation of the efficient Bayesian approach, as Zarifzadeh et al. (2024) empirically demonstrates that these approaches exhibit similar performances. Let $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ denote, respectively, the mean and standard deviation of $t_{b}$ estimated from shadow models, where $a$ denotes which of $(x,y_{x})$ and $(z,y_{z})$ is in the training set. By Equation 11 in (Zarifzadeh et al., 2024), the per-example performance for RMIA is given as
$\textsc{tpr}_{\mathrm{RMIA}}(x)=\\ \Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Pr_{(z,y_{z})\sim{\mathbb{D}}}(\mathrm{LR}(x,z)\geq\gamma)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\right)$ (7)
$\mathrm{LR}(x,z)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{x,z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{z,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z,z}^{2})},$ (8)
where $t_{z}$ is the score on $z$ similar to $t_{x}$ and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Appendix A for derivation).
@@
We define the average-case tprs for LiRA and RMIA by taking the expectation over the data distribution:
$\overline{\textsc{tpr}}_{\mathrm{LiRA}}=\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(x)]$ (9)
$\overline{\textsc{tpr}}_{\mathrm{RMIA}}=\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(x)]$ (10)

### v4

@@ Abstract
Membership inference attacks (MIAs) are used to test practical privacy of machine learning models. MIAs complement formal guarantees from differential privacy (DP) under a more realistic adversary model. We analyse MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning. We show that the vulnerability of non-DP models when measured as the attacker advantage at fixed false positive rate reduces according to a simple power law as the number of examples per class increases, even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
@@ 1 Introduction
Membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and differential privacy (DP; Dwork et al., 2006) provide complementary means of deriving lower and upper bounds for the privacy loss of a machine learning algorithm. Yet, the two operate under slightly different threat models. DP implicitly assumes a very powerful adversary with access to all training data except the target point and provides guarantees against every target point. MIAs assume an often more realistic adversary model with access to just the data distribution, and the unknown training data becoming latent variables introducing stochasticity into the attack, but the practical evaluation is statistical and cannot provide universal guarantees.
In this paper, we seek to explore the area between DP and typical MIA evaluations. Inspired by an empirical finding that average MIA vulnerability of neural network fine-tuning strongly reduces as the number of samples in the target class increases (see Figure 1), we develop theory of optimal MIA against a simple model of neural network fine-tuning and reproduce the decrease in vulnerability. Furthermore, the theoretical model predicts that the vulnerability of all individual samples should reduce as the number of samples increases, which we are able to verify empirically.
Figure 1: We observe a power-law relation between MIA vulnerability and $S$ (shots, i.e., examples per class) when attacking a ViT-B Head fine-tuned using LiRA. The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over six seeds.
To achieve our goal, we theoretically analyze and systematically apply two state-of-the-art black-box MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections. For the theoretical model that we analyse, LiRA is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933). Under the black-box threat model, in which the adversary does not have access to the model parameters, LiRA and RMIA have been shown to empirically outperform other attacks, especially when the number of shadow models is sufficiently large.
We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning and especially important when labeled examples are limited, which is often the case in privacy-sensitive applications. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA. In particular, we theoretically study the relationship between the number of examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning, and derive a power-law relationship (Figure 1) in the form
@@
We complement the theoretical analysis with extensive experiments over many datasets with varying sizes, in the transfer learning setting for image classification tasks, and observe the same power-law. Based on extrapolation from our results, the number of examples per each class that are needed for adequate protection of the most vulnerable samples appears very high.
Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020). A larger generalisation error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018). Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., ; Bagdasaryan et al., 2019). Feldman and Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning. Additionally, Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots. Yu et al. (2023) studied the relationship between the MIA vulnerability and individual privacy parameters for different classes. Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2024). Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change. Our work significantly expands on these works by explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation 1.
List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art score-based MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), which are a strong realistic threat model. We first analytically derive the power-law relationship in Equation 1 for both MIAs by introducing a simplified model of the optimal membership inference (Section 3). We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models by focusing on a transfer-learning setting for image classification task, where a large pre-trained neural network is fine-tuned on a sensitive dataset.
@@
    Power-law in a simplified model of the optimal MIA: We formulate a simplified model of MIA to quantitatively relate dataset properties and MIA vulnerability. In this model LiRA is the optimal attack. For this model, we prove a power-law relationship for both average and worst-case between the LiRA as well as RMIA vulnerability and the number of examples per class (See Section 3.4).
@@
    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures 1 and 4.2). We fit a regression model which follows the functional form of the theoretically derived power-law. We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section 4.3).
@@
    MIA experiments on the worst-case vulnerability: We extend the experiments to worst-case individual sample vulnerabilities and observe a similar decrease in vulnerability for quantiles of vulnerable data points and a slower decrease for the maximum individual vulnerability (Section 4.4). By extrapolation we find that an adequate protection of the most vulnerable samples would require an extremely large dataset (Section 4.5).
@@ 2 Background
Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model (Shokri et al., 2017). Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through DP.
@@
LiRA exploits the observation that the loss function value used to train a model is often lower for the examples that were part of the training set compared to those that were not. For a target tuple $(\bm{x},y)$, where $y$ is a label, LiRA trains the shadow models: 1 with $(\bm{x},y)$ as a part of the training set ($(\bm{x},y)\in\mathcal{D}$) and 2 without $\bm{x}$ in the training set ($(\bm{x},y)\notin\mathcal{D}$) . After training the shadow models, $(\bm{x},y)$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $(\bm{x},y)\in\mathcal{D}$ shadow models, and one for the $(\bm{x},y)\notin\mathcal{D}$. Finally, the attacker computes the loss for the point $\bm{x}$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $(\bm{x},y)\in\mathcal{D}$ or $(\bm{x},y)\notin\mathcal{D}$. When the true distributions of the shadow models are Gaussians, LiRA is the optimal attack provided by the Neyman–Pearson lemma (Neyman and Pearson, 1933). We use an optimization by Carlini et al. (2022) for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models. It relies on sampling the shadow datasets in a way that each sample is in expectation half of the time included in the training dataset of a shadow model and half of the time not. At attack time each model will be attacked once using all other models as shadow models.
Robust Membership Inference Attack (RMIA; Zarifzadeh et al., 2024) RMIA is a new MIA algorithm, which aims to improve performance when the number of shadow models is limited. We show both theoretically and empirically that the power-law also holds for RMIA.
@@
Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples. Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators. Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals (Clopper and Pearson, 1934). Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
$\begin{split}&B(\alpha/2;\textsc{tp},\textsc{p}-\textsc{tp}+1)<\textsc{tpr}<B(1-\alpha/2;\textsc{tp}+1,\textsc{p}-\textsc{tp}),\end{split}$ (2)
where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
@@ 3 Theoretical analysis
In this section, we seek to theoretically understand the impact of the dataset properties on the MIA vulnerability. It is known that different data points exhibit different levels of MIA vulnerability depending on the underlying distribution (e.g.  Aerni et al., 2024; Leemann et al., 2024). Therefore, we start with analysing per-example MIA vulnerabilities. In order to quantitatively relate dataset properties to these vulnerabilities, a simplified model is formulated. Within this model, we prove a power-law between the per-example vulnerability and the number $S$ of examples per class. Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section 4. We primarily focus on the analysis of LiRA, since it is the optimal attack in our simplified model. In Appendix B we show that similar theoretical results also hold for RMIA.
@@ 3.1 Preliminaries
First, let us restate the MIA score from LiRA as defined by Carlini et al. (2022). Denoting the logit of a target model $\mathcal{M}$ applied on a target data point $(\bm{x},y)$ as $\ell(\mathcal{M}(\bm{x}),y)$, LiRA computes the MIA score as the likelihood ratio
$\mathrm{LR}(\bm{x})=\frac{p(\ell(\mathcal{M}(\bm{x}),y)\mid\mathbb{Q}_{\text{in}}(\bm{x},y))}{p(\ell(\mathcal{M}(\bm{x}),y)\mid\mathbb{Q}_{\text{out}}(\bm{x},y))},$ (3)
@@
where the $\mathbb{Q}_{\text{in}/\text{out}}$ denote the hypotheses that $(\bm{x},y)$ was or was not in the training set of $\mathcal{M}$. Carlini et al. (2022) approximate the IN/OUT hypotheses as normal distributions. Denoting $t_{\bm{x}}=\ell(\mathcal{M}(\bm{x}),y)$, the score becomes
$\mathrm{LR}(\bm{x})=\frac{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\sigma}_{\mathrm{in}}(\bm{x})^{2})}{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{out}}(\bm{x}),\hat{\sigma}_{\mathrm{out}}(\bm{x})^{2})},$ (4)
@@
where the $\hat{\mu}_{\text{in/out}}(\bm{x})$ and $\hat{\sigma}_{\text{in/out}}(\bm{x})$ are the means and standard deviations for the IN/OUT shadow model losses for $(\bm{x},y)$. Larger values of $\mathrm{LR}(\bm{x})$ suggest that $(\bm{x},y)$ is more likely in the training set and vice versa. Now, to build a classifier from this score, the LiRA tests if $\mathrm{LR}(\bm{x})>\tau$ for some threshold $\tau$. Note that the attacker only has a finite set of shadow models to estimate the IN/OUT parameters. Therefore, the MIA scores become random variables over the true population level IN/OUT distributions.
3.2 Computing the tpr for LiRA
Using the LiRA formulation of Equation 4, the tpr for the target point $(\bm{x},y)$ for LiRA is defined as
@@ 3.2 Computing the tpr for LiRA
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\mathrm{LR}(\bm{x})\geq\tau\mid(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\right),$ (5)
@@
where $\tau$ is a threshold that defines a rejection region of the likelihood ratio test, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Section A.1 for derivation).
@@
We define the average-case tpr for LiRA by taking the expectation over the data distribution:
$\overline{\textsc{tpr}}_{\mathrm{LiRA}}=\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})]$ (6)